OpenStudy (anonymous):

2y-6 -------- <---- DIVISION.SOMEONE PLEASE HELP 2y^2+5y+2

OpenStudy (anonymous):

That pellet don't help !

OpenStudy (anonymous):

pellet

OpenStudy (anonymous):

pellet?

OpenStudy (anonymous):

The page has an example that goes through this exact type of problem.

OpenStudy (anonymous):

sh It? What?

OpenStudy (anonymous):

put it together.

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

There is the link providing you with the information so that you can do the question...

OpenStudy (anonymous):

First job: follow along the example on that link, seeing how it works.

OpenStudy (anonymous):

doesn't really help, can you help me ?

OpenStudy (anonymous):

Tell me how the link doesn't help, and I will help you to follow the example in the link.

OpenStudy (anonymous):

because in my book its different, you have to factor the numerator and denominator, then use fundamental property, I really don't get hoe that works.

OpenStudy (anonymous):

how*

OpenStudy (anonymous):

Oh right, well I don't understand what method your book wants you to use.

OpenStudy (anonymous):

well the title is " The Fundamental Property of Rational Expressions " ?

OpenStudy (anonymous):

Well, good luck with that. I have to go.

OpenStudy (anonymous):

flutter you

OpenStudy (anonymous):

Nice way to talk to someone who voluntarily tried to help you.

OpenStudy (anonymous):

YOU DIDNT EVEN HELPL ME , YOUR WORTHLESS

OpenStudy (accessdenied):

There is simply no reason to bash others here. If you do not think that they were helpful, then tell them straight up: "sorry, that did not help me." There is no reason to engage caps lock mode and bring down the hate for somebody who volunteered just to help you, though. :s

OpenStudy (anonymous):

CAN YOU HELP ME ?

OpenStudy (accessdenied):

Maybe if you can stop the caps lock only messages. It's harder to reply if you're going to type like that... ;)

OpenStudy (anonymous):

@Andres909z the answers would be something like that: a+bx+cx^2+......

OpenStudy (anonymous):

sorry . but can you ?

OpenStudy (anonymous):

\[\frac{2y-6}{2y^2+5y+2}\] You can apply Horner

OpenStudy (anonymous):

horner?

OpenStudy (anonymous):

with horner you'll get a serie

OpenStudy (anonymous):

whats that

OpenStudy (accessdenied):

It doesn't appear that we have many options here for division. The numerator is smaller than the denominator. The only simplification we could do is maybe cancel like-factors in the numerator and denominator by factoring them and cancelling. After that, we'd have nothing else to do really.

OpenStudy (anonymous):

so what would be the answer to this problem ?

OpenStudy (anonymous):

-3 over 10y perhaps?

OpenStudy (accessdenied):

Well, when I factor: 2y - 6 = 2(y - 3) 2y^2 + 5y + 2 = 2y^2 + y + 4y + 2 = y(2y + 1) + 2(2y + 1) = (y + 2)(2y + 1) I can't see anything that cancels between the numerator and denominator. So it seems like there is nothing else to do here, we basically have the simplest form already...

OpenStudy (anonymous):

so what do I write as my final answer?

OpenStudy (accessdenied):

Something along the lines of that result -- it's already in simplest form, we cannot do anything else to this expression to simplify it...